# Position: Many generalization measures for deep learning are fragile

### v2

@@ A.1 Measures considered
We adopt the categories and normalizations used by prior experimental studies of generalization measures (Jiang et al., 2019b; Dziugaite et al., 2020b); for exact constants and implementation choices, see App. C.6 of Dziugaite et al. (2020b). Let a $d$-layer network have layer weights $\{W_{i}\}_{i=1}^{d}$ with initialization $\{W_{i}^{0}\}$; write $\|\cdot\|_{F}$ and $\|\cdot\|_{2}$ for Frobenius and spectral norms; let $n$ denote training-set size and $\gamma$ a robust (e.g., 10th-percentile) training margin.
@@
  • Combined spectral–Frobenius ratio. We follow the combined ratio used in prior large‑scale studies (App. C.6 of Dziugaite et al., 2020a), denoted FRO_OVER_SPEC, which normalizes Frobenius quantities by spectral ones to reduce raw scale effects. We report it alongside its constituents in our plots.
@@ A.6 VC-dimension proxy (robust baseline)
Across all families above, the figures make a consistent point: changing only the learning rate or swapping the optimizer can reshape the surrogate’s learning curve—introducing plateaus, rebounds, late spikes, or crossings—without a commensurate shift in test accuracy. This shape-level fragility extends far beyond Path norms and cautions against reading causal explanations of generalization from any single surrogate without dedicated stress tests (Dziugaite et al., 2020b; Jiang et al., 2019b).
@@ Appendix B Additional temporal behavior results: optimizer sensitivity across measure families
Appendix C Proof for theorem 6.2
In this section we prove Theorem 6.2. We first reintroduce the scale invariance lemma from Arora et al. (2018b); Li & Arora (2019) which is the key source of intuition about scale-invariant networks.
@@
We give the technical definition of the commonly used training algorithm SGD with momentum and weight decay (WD) (with respect to the L2 norm of the parameters) using a convenient form given in Li & Arora (2019), which is equivalent to the default implementation in Pytorch (Paszke et al., 2017).
@@
The range of $\alpha$ can be easily shown by combining the following two constraints and assuming remark 6.3 is true:
@@
Now we are ready to prove theorem 6.2.
Proof of Theorem 6.2.
@@ Appendix D Additional label-corruption results: PAC–Bayes and Path Norms
Appendix E Quantitative fragility: detailed results
This appendix extends Section 7. We first summarize the experimental setup, and then give concise, implementable procedures for our two instability scores: CMS and its seed‑adjusted variant eCMS. We close with a brief reading guide that highlights what can be learned from the summary tables at Table 1 ($\delta=0.01$), Table 2 ($\delta=0.02$), and Table 3 ($\delta=0.05$). For full definitions of the measures listed in those tables, see (Dziugaite et al., 2020b).
@@ Appendix E Quantitative fragility: detailed results
We study two datasets—FashionMNIST and CIFAR‑10—across three architectures: Network‑in‑Network (NiN) (Lin et al., 2013), ResNet‑50 (RN50), and DenseNet‑121 (DN121). The training pipeline varies along three axes: (i) learning rate $\texttt{lr}\in\{0.001,0.0032,0.0063,0.01,0.0158,0.05,0.1\}$; (ii) optimizer $\in\{\textsc{Adam},\textsc{SGD}+\textsc{momentum}\}$; and (iii) stopping criterion: either the first epoch reaching $100\%$ training accuracy, or the first epoch whose training cross‑entropy drops below $0.01$. Each hyperparameter configuration is replicated with $8$ random seeds. We compute scores within each (dataset, architecture) group at absolute error tolerances $\delta\in\{0.01,0.02,0.05\}$, using natural logs and reporting medians.
@@
We contrast post-training (“post-mortem”) generalization measures with a function-space approach based on the marginal-likelihood PAC–Bayes bound of Valle-Pérez & Louis (2020). The bound controls the test error of a hypothesis sampled from a Bayesian posterior over functions (not parameters); the key capacity term is the marginal likelihood (Bayesian evidence) of the data under the architecture’s Gaussian‑process (GP) limit. All figure numbers below refer to Valle-Pérez & Louis (2020).
@@
  • Data complexity. The marginal‑likelihood bound tracks dataset difficulty: it increases with label corruption and preserves the canonical ordering MNIST << Fashion‑MNIST << CIFAR‑10; see Figure 1 in (Valle-Pérez & Louis, 2020). In our audits, several post‑mortem measures—including PACBAYES_orig—either flattened under corruption or were not able to catch the ordering, even when the test error did.
  • Learning‑curve scaling. The marginal‑likelihood bound tracks learning‑curve scaling: it mirrors the empirical power‑law in $n$ and clusters exponents primarily by dataset; see Figures 2–5 in (Valle-Pérez & Louis, 2020). By contrast, some post‑mortem measures barely moved with $n$; others bent the wrong way or mixed optimizer effects with data effects (cf. (Nagarajan & Kolter, 2019)).
  • Temporal behavior (overtraining invariance). The marginal‑likelihood bound tracks temporal invariance: once $S$ is fixed and training has interpolated ($\hat{\varepsilon}=0$), the bound depends only on $P(C(S))$ and nn—it is agnostic to how long or by which path the parameters were trained. In our temporal‑behavior experiments, several norm‑based post‑mortem measures (e.g., $\ell_{2}$/spectral/path norms) kept growing during overtraining while the generalization error stayed roughly unchanged; the GP bound remained stable. (This echoes broader observations that longer training can leave test error flat or improved (Hoffer et al., 2017), and that norms can diverge under separable losses without hurting classification error (Soudry et al., 2018).)
  • Training‑pipeline changes. Many post‑mortem measures swing with optimizer, batch size, explicit/implicit regularization, and early stopping—sometimes tracking curvature artifacts rather than out‑of‑sample error (e.g., Keskar et al., 2017; Hochreiter & Schmidhuber, 1997; Jiang et al., 2019b; Dziugaite et al., 2020b). The GP marginal‑likelihood bound is, by construction, insensitive to these knobs: it depends on the architecture’s function prior and the data, not on the path taken through parameter space. We view this invariance as a virtue for a predictor; the bound is not designed to explain differences caused by optimizer choice and should not be judged on that criterion.
@@
Two ingredients stand out. First, the bound evaluates an Occam factor in function space: architectures that place high prior mass on data‑compatible functions earn better evidence, naturally capturing dataset ordering and learning‑curve slopes. Second, there is a credible external yardstick: infinite‑width GP predictors (NNGP/NTK) often quantitatively predict finite‑width DNN generalization trends (Lee et al., 2020a). Together, these explain why Figures 1–5 show greater stability for the GP‑based approach than we observe for post‑hoc, parameter‑space measures.
@@
The live hypothesis is that they measure the wrong object. Post‑mortem scores probe properties of a single trained parameter vector (curvature, norms, compressibility), entangling optimization details and reparameterization choices with generalization. That enterprise is important—and the community has argued forcefully that such post‑training diagnostics deserve attention (e.g., compression and robustness perspectives in Arora et al., 2018a, 2019; Jiang et al., 2019b; Dziugaite et al., 2020b)—but our evidence suggests their predictions are fragile under routine perturbations. By contrast, the GP bound targets the distribution over functions implied by architecture and data.
@@ Appendix H Exp++ in scale-invariant nets: protocol and results
We instantiate the symmetry in a fully connected, scale‑invariant network (normalization after each hidden linear layer; the final linear readout is frozen), trained on MNIST with SGD + momentum. We sweep a single control, the Exp++ multiplicative factor $\alpha$ from Theorem 6.2, which makes the learning rate grow exponentially across steps, and repeat the sweep both without and with weight decay. For each run we log parameter‑space PAC–Bayes proxies (including magnitude‑aware variants), path‑norm proxies (with/without margin normalization), and test error. The stopping rule is cross‑entropy unless stated otherwise.

### v3

@@ A.1 Measures considered
We adopt the categories and normalizations used by prior experimental studies of generalization measures (jiang2019fantastic; dziugaite2020in); for exact constants and implementation choices, see App. C.6 of dziugaite2020in. Let a $d$-layer network have layer weights $\{W_{i}\}_{i=1}^{d}$ with initialization $\{W_{i}^{0}\}$; write $\|\cdot\|_{F}$ and $\|\cdot\|_{2}$ for Frobenius and spectral norms; let $n$ denote training-set size and $\gamma$ a robust (e.g., 10th-percentile) training margin.
@@
  • Combined spectral–Frobenius ratio. We follow the combined ratio used in prior large‑scale studies (App. C.6 of dziugaite2020in), denoted FRO_OVER_SPEC, which normalizes Frobenius quantities by spectral ones to reduce raw scale effects. We report it alongside its constituents in our plots.
@@ A.6 VC-dimension proxy (robust baseline)
Across all families above, the figures make a consistent point: changing only the learning rate or swapping the optimizer can reshape the surrogate’s learning curve—introducing plateaus, rebounds, late spikes, or crossings—without a commensurate shift in test accuracy. This shape-level fragility extends far beyond Path norms and cautions against reading causal explanations of generalization from any single surrogate without dedicated stress tests (dziugaite2020in; jiang2019fantastic).
@@ Appendix B Additional temporal behavior results: optimizer sensitivity across measure families
Appendix C Theory of scale-invariance: definitions and proofs
In this section, we provide the formal definition of scale-invariant networks and the theorem regarding the equivalence of schedules (the ”Exp++” factor) discussed in Section 7. We then provide the full proof.
C.1 Definitions and Theorem Statement
Definition C.1 (Scale invariant neural networks).
Theorem C.2 (Equivalence of schedules in scale-invariant nets).
[…]
Define the interval endpoints
$\alpha_{L}$ $\coloneqq\frac{\gamma}{\,1-\lambda\eta_{0}+\gamma\,}$, (4)
and set
$\mathcal{I}\coloneqq(\alpha_{L},\ \alpha_{-}]\ \cup\ [\alpha_{+},\,1).$ (7)
If $\alpha\in\mathcal{I}$, then the schedules
$\tilde{\eta}_{t}$ $=\eta_{0}\,\alpha^{-2t-1},$ (8)
Remark C.3 (Parameter range and nonnegativity).
Assume
$\lambda\eta_{0}$ $\leq(1-\sqrt{\gamma})^{2},$ (10)
C.2 Proof of Theorem C.2
We first reintroduce the scale invariance lemma from arora2018theoretical; li2019an which is the key source of intuition about scale-invariant networks.
@@ C.1 Definitions and Theorem Statement
We give the technical definition of the commonly used training algorithm SGD with momentum and weight decay (WD) (with respect to the L2 norm of the parameters) using a convenient form given in li2019an, which is equivalent to the default implementation in Pytorch (paszke2017automatic).
@@
The range of $\alpha$ can be easily shown by combining the following two constraints and assuming remark C.3 is true:
@@
Now we are ready to prove theorem C.2.
Proof of Theorem C.2.
@@ Appendix D Additional label-corruption results: PAC–Bayes and Path Norms
Appendix E Quantitative fragility: detailed definitions and results
This appendix provides the formal definitions, algorithmic details, and full results for the quantitative analysis summarized in Section 8.
E.1 Formal Definitions
To quantify fragility, we define metrics that capture how much a generalization measure fluctuates among runs with nearly identical test error.
Setup.
Definition E.1 (Conditional Measure Spread (CMS)).
The CMS of $C$ in group $g$ at tolerance $\delta$ is
Randomness adjustment.
Define the corresponding spreads
Definition E.2 (eCMS (Excess CMS)).
The seed‑adjusted excess at matched error is
Aggregation across groups.
How to read the scores.
Properties and protocol.
Both statistics are robust (medians) and scale‑free (invariant under $C\mapsto aC$, $a>0$). Unless noted otherwise we use absolute tolerances $\delta\in\{0.01,\,0.02,\,0.05\}$ (i.e., $1\%$ - $5\%$ of error), uniformly subsample pairs if they exceed a fixed budget, and report, for each measure, the medians $\mathrm{CMS}_{\delta}^{\mathrm{med}}$ and $\mathrm{eCMS}_{\delta}^{\mathrm{med}}$, along with the number of eligible pairs in (15)–(18) and the fraction of groups covered. Together, CMS and eCMS capture (i) local instability at matched error and (ii) the seed‑adjusted component attributable to hyperparameters.
E.2 Experimental Details
We first summarize the experimental setup, and then give concise, implementable procedures for our two instability scores: CMS and its seed‑adjusted variant eCMS. We close with a brief reading guide that highlights what can be learned from the summary tables at Table 1 ($\delta=0.01$), Table 2 ($\delta=0.02$), and Table 3 ($\delta=0.05$). For full definitions of the measures listed in those tables, see (dziugaite2020in).
@@ E.2 Experimental Details
We study two datasets—FashionMNIST and CIFAR‑10—across three architectures: Network‑in‑Network (NiN) (lin2013network), ResNet‑50 (RN50), and DenseNet‑121 (DN121). The training pipeline varies along three axes: (i) learning rate $\texttt{lr}\in\{0.001,0.0032,0.0063,0.01,0.0158,0.05,0.1\}$; (ii) optimizer $\in\{\textsc{Adam},\textsc{SGD}+\textsc{momentum}\}$; and (iii) stopping criterion: either the first epoch reaching $100\%$ training accuracy, or the first epoch whose training cross‑entropy drops below $0.01$. Each hyperparameter configuration is replicated with $8$ random seeds. We compute scores within each (dataset, architecture) group at absolute error tolerances $\delta\in\{0.01,0.02,0.05\}$, using natural logs and reporting medians.
@@
We contrast post-training (“post-mortem”) generalization measures with a function-space approach based on the marginal-likelihood PAC–Bayes bound of valleprez2020generalization. The bound controls the test error of a hypothesis sampled from a Bayesian posterior over functions (not parameters); the key capacity term is the marginal likelihood (Bayesian evidence) of the data under the architecture’s Gaussian‑process (GP) limit. All figure numbers below refer to valleprez2020generalization.
@@
  • Data complexity. The marginal‑likelihood bound tracks dataset difficulty: it increases with label corruption and preserves the canonical ordering MNIST << Fashion‑MNIST << CIFAR‑10; see Figure 1 in (valleprez2020generalization). In our audits, several post‑mortem measures—including PACBAYES_orig—either flattened under corruption or were not able to catch the ordering, even when the test error did.
  • Learning‑curve scaling. The marginal‑likelihood bound tracks learning‑curve scaling: it mirrors the empirical power‑law in $n$ and clusters exponents primarily by dataset; see Figures 2–5 in (valleprez2020generalization). By contrast, some post‑mortem measures barely moved with $n$; others bent the wrong way or mixed optimizer effects with data effects (cf. (nagarajan2019uniform)).
  • Temporal behavior (overtraining invariance). The marginal‑likelihood bound tracks temporal invariance: once $S$ is fixed and training has interpolated ($\hat{\varepsilon}=0$), the bound depends only on $P(C(S))$ and nn—it is agnostic to how long or by which path the parameters were trained. In our temporal‑behavior experiments, several norm‑based post‑mortem measures (e.g., $\ell_{2}$/spectral/path norms) kept growing during overtraining while the generalization error stayed roughly unchanged; the GP bound remained stable. (This echoes broader observations that longer training can leave test error flat or improved (hoffer2017train), and that norms can diverge under separable losses without hurting classification error (soudry2018the).)
  • Training‑pipeline changes. Many post‑mortem measures swing with optimizer, batch size, explicit/implicit regularization, and early stopping—sometimes tracking curvature artifacts rather than out‑of‑sample error (e.g., keskar2016on; hochreiter1997flatb; jiang2019fantastic; dziugaite2020in). The GP marginal‑likelihood bound is, by construction, insensitive to these knobs: it depends on the architecture’s function prior and the data, not on the path taken through parameter space. We view this invariance as a virtue for a predictor; the bound is not designed to explain differences caused by optimizer choice and should not be judged on that criterion.
@@
Two ingredients stand out. First, the bound evaluates an Occam factor in function space: architectures that place high prior mass on data‑compatible functions earn better evidence, naturally capturing dataset ordering and learning‑curve slopes. Second, there is a credible external yardstick: infinite‑width GP predictors (NNGP/NTK) often quantitatively predict finite‑width DNN generalization trends (lee2020finite). Together, these explain why Figures 1–5 show greater stability for the GP‑based approach than we observe for post‑hoc, parameter‑space measures.
@@
The live hypothesis is that they measure the wrong object. Post‑mortem scores probe properties of a single trained parameter vector (curvature, norms, compressibility), entangling optimization details and reparameterization choices with generalization. That enterprise is important—and the community has argued forcefully that such post‑training diagnostics deserve attention (e.g., compression and robustness perspectives in arora2018strongerc; arora2019fine; jiang2019fantastic; dziugaite2020in)—but our evidence suggests their predictions are fragile under routine perturbations. By contrast, the GP bound targets the distribution over functions implied by architecture and data.
@@ Appendix H Exp++ in scale-invariant nets: protocol and results
We instantiate the symmetry in a fully connected, scale‑invariant network (normalization after each hidden linear layer; the final linear readout is frozen), trained on MNIST with SGD + momentum. We sweep a single control, the Exp++ multiplicative factor $\alpha$ from Theorem C.2, which makes the learning rate grow exponentially across steps, and repeat the sweep both without and with weight decay. For each run we log parameter‑space PAC–Bayes proxies (including magnitude‑aware variants), path‑norm proxies (with/without margin normalization), and test error. The stopping rule is cross‑entropy unless stated otherwise.
@@
Appendix I Additional Compression Bounds Results
In Section 6, we analyzed the breakdown of compression bounds for CIFAR-10 (Intrinsic Dimension 3500) and MNIST. Here, we provide the corresponding omitted results for CIFAR-10 with a higher intrinsic dimension ($ID=10000$).
(a) CIFAR-10 (ID 10000) Bounds
(b) CIFAR-10 (ID 10000) Raw KL
Figure S20: Compression Bounds with Higher Intrinsic Dimension (CIFAR-10). These plots correspond to the setting where the intrinsic dimension is set to 10,000. Left: The bound components versus sample size. Right: The raw KL divergence. Consistent with the main text analysis, the complexity term (KL) remains flat despite the increase in sample size.